OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP!! A(n) 866 N crate is being pushed across a level floor by a force of 350 N at an angle of 21◦ above the horizontal. The coefficient of kinetic friction between the crate and the floor is 0.24.  The acceleration of gravity is 9.81 m/s2 What is the acceleration of the box? Answer in units of m/s2? I keep getting 1.688045635 as the answer! BUT ITS WRONG! If you get a different answer, please show me how and give an exact number such as mine, thanks a bunch!

OpenStudy (maheshmeghwal9):

is 866N crate's weight?

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

ok

OpenStudy (shubhamsrg):



OpenStudy (maheshmeghwal9):

Is 1.0239m/s2 is the answer? @natasha.aries

OpenStudy (shubhamsrg):

N = 866 - 350 sin21 350 cos21 - 0.24 N = (866/9.8) a use both to solve for a

OpenStudy (anonymous):

is that what both of you got?

OpenStudy (shubhamsrg):

i didnt calculate,,am too lazy doing so ! :P

OpenStudy (maheshmeghwal9):

ya i gt \[1.0239m/s^2.\] @natasha.aries

OpenStudy (anonymous):

no that it not right :/

OpenStudy (shubhamsrg):

i get 1.68632 (used wolfram)

OpenStudy (maheshmeghwal9):

then i think @shubhamsrg would do it:)

OpenStudy (anonymous):

as the whole answer?

OpenStudy (anonymous):

no that is incorrect as well

OpenStudy (maheshmeghwal9):

\[1m/s^2 \space or \space 2m/s^2\]

OpenStudy (maheshmeghwal9):

as an approximate

OpenStudy (maheshmeghwal9):

wt is the answer @natasha.aries ?

OpenStudy (anonymous):

idk i have to submit it online and it just tells me if its incorrect or right

OpenStudy (maheshmeghwal9):

oh i see-_-

OpenStudy (anonymous):

yeah :/

OpenStudy (amriju):

ok....u need to draw the free body diagram...as its calld over here ,first nowu can understand that the horizontal component of ur force is responsible for the accelaration but there is friction too....now horizontal comp of force...350cos21 thats 326.753 N. now then the fric force....we know that it depends on the normal...and here the normal force is not the weight but less than it...since a part of ur applied force is acting against the gravitational force thats 350sin21=125.428 N. then net normal force= 866-125.428=740.572 N....fric=740.572*.24=177.737. now then net force= 326.753-177.737=149.016 mass of block=866/9.81=88.277 k accn=1.688